# A Survey on Automatic Credibility Assessment of Textual Credibility Signals in the Era of Large Language Models

### v1

@@ Abstract.
In the current era of social media and generative AI, an ability to automatically assess the credibility of online social media content is of tremendous importance. Credibility assessment is fundamentally based on aggregating credibility signals, which refer to small units of information, such as content factuality, bias, or a presence of persuasion techniques, into an overall credibility score. Credibility signals provide a more granular, more easily explainable and widely utilizable information in contrast to currently predominant fake news detection, which utilizes various (mostly latent) features. A growing body of research on automatic credibility assessment and detection of credibility signals can be characterized as highly fragmented and lacking mutual interconnections. This issue is even more prominent due to a lack of an up-to-date overview of research works on automatic credibility assessment. In this survey, we provide such systematic and comprehensive literature review of 175 research papers while focusing on textual credibility signals and Natural Language Processing (NLP), which undergoes a significant advancement due to Large Language Models (LLMs). While positioning the NLP research into the context of other multidisciplinary research works, we tackle with approaches for credibility assessment as well as with 9 categories of credibility signals (we provide a thorough analysis for 3 of them, namely: 1) factuality, subjectivity and bias, 2) persuasion techniques and logical fallacies, and 3) claims and veracity). Following the description of the existing methods, datasets and tools, we identify future challenges and opportunities, while paying a specific attention to recent rapid development of generative AI.
@@ 1. Introduction
Tackling false information (i.e., disinformation and misinformation) has attracted significant attention in recent years from researchers, media professionals, AI/social media practitioners as well as a general public. From the research perspective, a considerable focus emerged notably after the US presidential elections in 2016. Since then, researchers have been addressing false information by means of a wide range of classification task, leveraging AI technologies such as machine learning, natural language processing, computer vision, or social network analysis (Sharma et al., 2019). The primary stream of research works, commonly denoted as a fake news detection, focuses on predicting whether a piece of content (considering different modalities and context) contains false information111Some works go beyond a binary classification and predict multiple classes, such as veracity labels with varying degrees of trustworthiness commonly used by fact-checking organizations: true, mostly true, mixture, mostly false, false. (Shu et al., 2017). Proposed solutions to fake news detection can be broadly categorized into style-based, propagation-based, source-based, and knowledge-based (Zhou and Zafarani, 2020). Style-based methods analyze linguistic and stylistic features of content, propagation-based methods focus on the spread patterns within networks, and source-based approaches assess the credibility of the information’s origin. Knowledge-based methods, on the other hand, involve content verification (fact-checking) against reliable sources, such as research studies, encyclopedias, knowledge graphs, or even large language models.
Despite commendable efforts with a clear contribution to protecting societal democratic values, proposed solutions for fake news detection have been recognized to lack (i) sufficient accuracy on out-of-distribution data (mostly due to over- or under-fitting of the detection models) (Hamed et al., 2023), (ii) generalizability (to new topics and emerging techniques for creating and spreading false information) (Suprem et al., 2022) and (iii) explainability (especially in the case of state-of-the-art deep learning models) (A.b. et al., 2023). All these factors play a crucial role in effectively combating false information.
Since these drawbacks of fake news detection are inherently related to the high-level nature of the task itself, a sibling task of credibility assessment has a potential to bring more accurate, better generalizable and explainable solutions as it is rather based on a two step approach. Firstly, a more granular credibility signals (sometimes denoted also as credibility indicators) are detected. Secondly, such credibility signals are aggregated by a credibility assessment algorithm resulting into a single credibility score. Typical examples of credibility signals include factuality, subjectivity and bias present in the content, persuasion techniques, logical fallacies or detection of recently emerging machine-generated text. While credibility signals can be analyzed in a piece of online content by a user manually, they provide a real value especially when detected automatically. Such automatic detection of credibility signals can range from simple (e.g., rule- or heuristic- based) approaches to complex solutions requiring AI that is capable of in-depth content understanding.
@@
As a result, credibility assessment provides an interesting potential to extend and supplement the existing stream of works on fake news detection. Credibility signals are more flexible and more widely-applicable as a binary/multiclass classification of fake news. Besides their primary usage, to be further aggregated into a credibility score, they can also be utilized and interpreted by a user directly (e.g., as an “information nutrition label” (Fuhr et al., 2018)), provide valuable inputs to information retrieval engines or recommender systems in order to prefer content/sources associated with a high level of credibility, or even serve as features in subsequent classification tasks (including fake news detection itself (Sitaula et al., 2020)). Study by Lu et al. (2022) showed that even if people are influenced by others when judging the veracity of online content, providing accurate AI-based credibility indicators can effectively improve people’s ability in detecting misinformation. Credibility assessment and credibility signals are, furthermore, especially useful in cases in which we cannot easily ascertain that something is true or false (on a single dimension) – concept of credibility provides a necessary level of granularity to represent a potentially complex information from multiple perspectives.
@@
Credibility signals are also more easily interpretable and naturally support semi-automatic human-centered AI approaches that attempt supporting instead of replacing a human expert or a lay person (individual credibility signals can be automatically detected and their interpretation can be subsequently done by a human). Last but not least, they are more aligned with journalistic/fact-checking workflows for identifying false information.
@@
In contrast to other related research areas (including fake news detection), research works on automatic credibility assessment are mostly carried out in isolation (as we show in this survey, there is a lack of works addressing multiple categories of credibility signals at once and, moreover, many research works do not explicitly state that the outcome of their solution can actually serve as a credibility signal). Getting familiar with the existing works on automatic credibility assessment and detection of credibility signals is challenging due to an ambiguous and inconsistent terminology used by researchers, a lack of clear definition of credibility and credibility signals as well as due to a missing standardized taxonomy of various credibility signals’ categories. At the same time, automatic credibility assessment can significantly benefit from a holistic approach. For example, due to common underlying similarities, there is a big potential of multi-task learning.
@@
While there is a wide range of possible categories of credibility signals, we specifically focus on such signals that relate to textual content and can be automatically detected by Natural Language Processing (NLP) methods. This area is currently at the spotlight due to the recent advancements of Large Language Models (LLMs). Their rapid development has a revolutionizing effect on many text classification tasks, automatic credibility assessment and detection of credibility signals not being an exception.
@@
Despite multiple surveys addressing fake news detection (e.g., (Hu et al., 2024; Tufchi et al., 2023; A.b. et al., 2023; Zhou and Zafarani, 2020; Shu et al., 2017)), to the best of our knowledge, there is no survey on automatic credibility assessment and detection of credibility signals from the NLP perspective. In this paper, we therefore provide the first comprehensive and systematic literature overview of credibility signal detection with the focus on NLP and LLMs. The main contributions are as follows:
@@
We provide a necessary background overview of definitions and dimensions of credibility assessment (including various taxonomies of credibility signals used in the existing literature) to setup the common understanding that is currently missing in the existing works.
@@
We analyze and describe a total of 175 NLP works tackling with automatic credibility assessment and detection of credibility signals. We provide an in-depth description of 3 major categories of credibility signals, which have been selected due to a considerable NLP research interest and their recognition as important by end users, namely: (i) factuality, subjectivity and bias; (ii) persuasion techniques and logical fallacies; and (iii) claims and veracity. This analysis is complemented with an overview of additional 6 categories of credibility signals.
We thoroughly position the NLP research on credibility assessment into the context of other modalities as well as multidisciplinary works. Gaps in the existing literature, future challenges and opportunities, considering also the rapid development of generative AI (especially LLMs), are discussed.
@@
Besides the review of published research works, this survey builds on extensive past research activities and acquired knowledge of its authors in the target area. A unique composition of authors consisting of experts not only on the primary NLP area, but also on other related research domains (such as a computer vision, a social network analysis) as well as media professionals, provides a novel holistic perspective on the addressed topic. Such diversity contributes, besides others, to identification of open problems and future challenges, in which an application of research in practice plays a prominent role.
@@
This survey is structured into 10 sections. Section 2 provides definitions of core concepts related to credibility, describes various dimensions of credibility assessment as well as defines the scope of this survey while situating it into the context of other existing surveys. In Section 3, the methodology employed in the survey process is thoroughly described. Section 4 focuses on research papers that address automatic credibility assessment (i.e., approaches addressing both steps — detection of multiple categories of credibility signals and their aggregation into a credibility score). Subsequently, Sections 5-7 address an automatic detection of 3 selected categories of credibility signals. Section 8 supplements this analysis with a brief overview of additional 6 categories of credibility signals, including the complementary perspective of credibility signals addressed in the non-NLP research. In Section 9, we provide an orthogonal analysis of challenges and open problems that characterise the state of the art in this area. Finally, conclusions are drawn in Section 10.
@@ 2.1. Definitions
Credibility is a complex and multidimensional concept which lacks a single, unified and widely-accepted definition. Existing definitions vary especially across various disciplines and fields, which study credibility from different perspectives, such as information science, psychology, marketing or human-computer interaction (HCI) (Rieh and Danielson, 2007). It is commonly defined as a high-level construct with a help of related concepts, such as believability (which is considered to be roughly a synonym with credibility), trust/trusthworthiness, reliability or their various combinations (Qureshi et al., 2021; Alrubaian et al., 2019; Rieh and Danielson, 2007; Self and Roberts, 2019). Credibility is always tied to a target entity (an object of assessment), that can be either a piece of content or a source spreading such content (Kiousis, 2001). Due to ambiguity of terminology, in this survey we opt for community-contributed definitions (Hawke, [n. d.]) gathered by the Credible Web Community Group222https://credweb.org/.
Credibility assessment is a process of ascertaining some level/degree of credibility to a target entity. Practically, credibility assessment can be formulated as (Qureshi et al., 2021): (i) a classification problem (i.e., assigning a label from a pre-defined categories, such as a low/high level of credibility); or as (ii) ranking/scoring problem (i.e., assigning a numerical score). In both cases, there are no standard scales adopted yet. Credible Web Community Group distinguishes two kinds of credibility assessment acts (Hawke, [n. d.]):
Making credibility decisions, the act of deciding for oneself what to believe, which is often informal, immediate, and unconscious. Doing this incorrectly often leads to being misled, although it may not be practical to do it correctly at all times.
Credibility analysis, manually by people or automatically by machines, is gathering, organizing, and analyzing evidence to help people make credibility decisions about a particular information item. A credibility analysis process might produce some kind of report which might itself be called a credibility assessment and might include a credibility score. In the following text, we will focus especially on this act.
@@
Credibility indicator (Hawke, [n. d.]) is commonly used interchangeably with credibility signal. In some communities, credibility signal is used for inputs to credibility assessment algorithms and credibility indicator is used for the display features added to the output, to communicate results to a human consumer. In this survey, we prefer to use credibility signal, while we would like to emphasize that in practice many signals can be used directly also as indicators communicated to end users.
Finally, credibility assessment tools (Hawke, [n. d.]) are software features or applications which perform credibility assessment or help people do so. Such tools can either derive credibility signals, or implement also a credibility assessment algorithm to aggregate such signals into a credibility score/credibility indicators.
Credibility and credibility assessment is inherently very close to false information (including fake news that intentionally create and spread disinformation) and the task of fake news detection. Nevertheless, credibility (as also defined above) is a broader concept and non-credible content does not necessary needs to be false one only. We would like to emphasize that credibility signals have been already recognized in fake news research as a possible solution towards more accurate and explainable methods. The recent work by Grieve and Woodfield (2023) delve into the complexities of identifying and analysing disinformation deceptive news. The authors study the way language adapts to different contexts and purposes, arguing that the linguistic choices in deceptive news differ systematically from those in genuine news. The difference in intent – to deceive versus to inform – leads to detectable variations in language use. The authors illustrate this with the famous case of fake news involving Jayson Blair of The New York Times, where false news were less informationally dense and less confident than real ones. The pressure to invent news and the lack of factual grounding influenced writing style, making it less precise and authoritative. The authors propose that this type of register variation could be a broader indicator of fake news. They call for continued research to understand the linguistic patterns of fake news and to develop strategies to combat its spread. Although the above framework is only based on the analysis of one reporter’s articles, this approach may be extended to corpora of non-credible/fake news and credible/real news on various topics to try to identify credibility signals/register variation between the two.
@@ 2.2. Dimensions of credibility assessment
Approaches to study credibility. First, the research works on credibility and credibility assessment can be broadly separated into multiple groups following their primary objective (Alrubaian et al., 2019).
There is a considerable number of (typically highly multidisciplinary) works focusing on definition of credibility, credibility components or theoretical frameworks.
The second group corresponds to human-based approaches (studies), in which various aspects of credibility are studied from the perspective of end-users.
Finally, a large group of approaches focuses on automation of credibility assessment, either detecting the credibility signals or aggregating them into the credibility indicators. The complexity of such automation can be very diverse, ranging from simple rule- and heuristic-based methods to supervised/unsupervised machine learning. This group also covers research works focusing on building datasets appropriate to create such methods/models; as well as systems (web applications, browser extensions, etc.) built on the top of such automated methods.
@@
Content level. At the content level, the task is to analyze the content attributes typically of a social media post or a news article. It is the most fundamental and important type of assessment, which is commonly used to determine the credibility at other levels (e.g., if a post is assessed to be credible, also a corresponding user/source or an associated topic/event is considered to be credible).
@@
First, many works broadly distinguish between content- and context-based credibility signals (Zhang et al., 2018; Kang et al., 2012; Bhuiyan et al., 2020; Wu et al., 2020; El Ballouli et al., 2017; Esteves et al., 2018; Qureshi and Malick, 2023; Piccolo et al., 2021; Jahanbakhsh and Karger, 2024; Chang et al., 2021; Ji et al., 2023b; Gao et al., 2015; Choi et al., 2023). Content-based signals are derived from the content itself, such as partisanship, emotional appeal, or persuasion techniques. On the other hand, context-based signals, also referred to as provenance- (Feng et al., 2023) or meta-information, take into account contextual (e.g., user, source) clues, such as user statistics, location, domain reputation, or the number of shares.
Second, despite many commendable efforts, there is no unified list or categorization of credibility signals. As a result, research works commonly utilize various and very diverse lists of credibility signals. For example, Molina et al. (2021) proposed separate lists of credibility signals (or features as they have been denoted by the authors) to distinguish between eight types of online content, such as real news, fabricated news, or satire. Each list grouped signals into 4 main categories roughly corresponding to levels of credibility assessment: (i) message and linguistic, (ii) sources and intentions, (ii) structural, and (iv) network.
An alternative approach was used by the Credibility Coalition333https://credibilitycoalition.org/, which organized weekly remote sessions during which participants drafted about 100 credibility signals specifically aimed at the credibility of web pages (Zhang et al., 2018). They have been later coalesced into 12 major categories, including reader behavior, revenue models, publication metadata, and inbound and outbound references. In 2017, the Credibility Coalition formed the World Wide Web Consortium (W3C) Credible Web Community Group, which continued in this initiative. It crowdsourced from human experts (researchers as well practitioners) an extensive list containing more than 200 credibility signals, commonly referred to as W3C signals (Multiple authors and Hawke, 2019). This list introduces a number of signals categorized by target entity (subject) type: claim, text, image, audio, video, article, title, web page, website, aggregation (e.g., RSS), venue, provider, creator, person, or organization. While such a list is currently just an informal draft and has not been standardized yet, it has been adopted and served as a foundation to several research works (e.g., (Piccolo et al., 2021; Podgurski et al., 2022)).

### v2

# A Survey on Automatic Credibility Assessment Using Textual Credibility Signals in the Era of Large Language Models

@@ Abstract.
In the age of social media and generative AI, the ability to automatically assess the credibility of online content has become increasingly critical, complementing traditional approaches to false information detection. Credibility assessment relies on aggregating diverse credibility signals – small units of information, such as content subjectivity, bias, or a presence of persuasion techniques – into a final credibility label/score. However, current research in automatic credibility assessment and credibility signals detection remains highly fragmented, with many signals studied in isolation and lacking integration. Notably, there is a scarcity of approaches that detect and aggregate multiple credibility signals simultaneously. These challenges are further exacerbated by the absence of a comprehensive and up-to-date overview of research works that connects these research efforts under a common framework and identifies shared trends, challenges, and open problems. In this survey, we address this gap by presenting a systematic and comprehensive literature review of 175 research papers, focusing on textual credibility signals within the field of Natural Language Processing (NLP), which undergoes a rapid transformation due to advancements in Large Language Models (LLMs). While positioning the NLP research into the the broader multidisciplinary landscape, we examine both automatic credibility assessment methods as well as the detection of nine categories of credibility signals. We provide an in-depth analysis of three key categories: 1) factuality, subjectivity and bias, 2) persuasion techniques and logical fallacies, and 3) check-worthy and fact-checked claims. In addition to summarising existing methods, datasets, and tools, we outline future research direction and emerging opportunities, with particular attention to evolving challenges posed by generative AI.
@@ 1. Introduction
Tackling false information (i.e., disinformation and misinformation) has attracted a substantial attention in recent years from researchers, media professionals, AI and social media practitioners as well as the general public. Researchers have approached false information detection through a variety of classification tasks, leveraging AI technologies such as machine learning, natural language processing, computer vision, or social network analysis (Sharma et al., 2019). The primary stream of research works on false information detection, commonly also denoted as a fake news detection (Shu et al., 2017; Zhou and Zafarani, 2020; Hu et al., 2024), assesses the veracity (i.e., factual accuracy) of information by means of a single-label prediction, typically a binary one (i.e., false/true content) (Shu et al., 2017). Some works go beyond a binary classification and predict multiple classes, such as veracity levels, commonly used by fact-checking organizations (e.g., true, mostly true, mixture, mostly false, false).
However, veracity alone offers only a limited perspective on the potential harm of online content. In practice, it should be considered alongside credibility. By proceeding from existing (quite heterogeneous) definitions (Hellmueller and Trilling, 2012; Qureshi et al., 2021; Alrubaian et al., 2019; Rieh and Danielson, 2007; Self and Roberts, 2019), under the credibility we understand the perceived trustworthiness, accuracy, and reliability of information or its source. It reflects the extent to which content is believed to be factual, unbiased, and free from manipulation or deception. False content that is perceived as highly credible by its audience can lead to significantly greater harm than false content that is clearly perceived as lacking credibility.
Credibility assessment therefore provides an interesting potential to extend and complement the existing stream of works on false information detection. It typically follows a two-step process. First, granular credibility signals (also referred to as credibility indicators) are detected. Second, these signals (serving as evidence) are aggregated into a single ordinal credibility label or a numerical credibility score. Common examples of credibility signals include subjectivity and bias present in the content, persuasion techniques, logical fallacies or presence of machine-generated content, which has become increasingly relevant with the rise of generative AI. While credibility assessment can be conducted manually by the end users, its full potential lies in automated detection of credibility signals and their consequent aggregation by a credibility assessment algorithm. These automated approaches may range from simple (e.g., rule-based or heuristic) methods to more advanced systems powered by AI, capable of nuanced and context-aware content analysis. Such automation not only improves scalability but also enables real-time evaluation of large volumes of online content.
@@
The benefit of credibility signals is their flexibility and wide-applicability. Besides their primary usage, to be aggregated into a credibility label/score, they can also be utilized and interpreted by a user directly (e.g., as an “information nutrition label” (Fuhr et al., 2018)), provide valuable inputs to information retrieval engines or recommender systems in order to prefer content/sources associated with a high level of credibility (Weerkamp and De Rijke, 2008), or even serve as features in subsequent classification tasks (including false information detection itself (Sitaula et al., 2020)). Credibility assessment and credibility signals are, furthermore, especially useful in cases in which we cannot easily ascertain that something is true or false (on a single veracity dimension) – concept of credibility provides a necessary level of granularity to represent a potentially complex information from multiple perspectives.
@@
Credibility signals are also more easily interpretable and naturally support semi-automatic human-centered AI approaches that attempt supporting instead of replacing a human expert or a lay person (individual credibility signals can be automatically detected and their interpretation can be subsequently done by a human). Study by Lu et al. (2022) showed that even if people are influenced by others when judging the veracity of online content, providing accurate AI-based credibility indicators can effectively improve people’s ability in detecting false information. Last but not least, they are also aligned with journalistic/fact-checking workflows for identifying trustworthy information (to be cited) or potential false information (to be fact-checked).
@@
In contrast to other related research areas (including false information detection), research works on automatic credibility assessment are mostly carried out in isolation. As we show in this survey, there is a lack of works addressing multiple categories of credibility signals at once and, moreover, many research works do not explicitly state that the outcome of their solution can actually serve as a credibility signal. Getting familiar with the existing works on automatic credibility assessment and detection of credibility signals is challenging due to an ambiguous and inconsistent terminology used by researchers, a lack of clear definition of credibility and credibility signals as well as due to a missing standardized taxonomy of various credibility signals’ categories. At the same time, automatic credibility assessment can significantly benefit from a holistic approach. For example, due to common underlying similarities, there is a big potential of transfer or multi-task learning.
@@
While there is a wide range of possible credibility signals, we specifically focus on such signals that relate to textual content and can be automatically detected by Natural Language Processing (NLP) methods. This area is currently at the spotlight due to the recent advancements of Large Language Models (LLMs). Their rapid development has a revolutionizing effect on many text classification tasks, automatic credibility assessment and detection of credibility signals not being an exception.
@@
Despite multiple surveys addressing false information detection (e.g., (Hu et al., 2024; Tufchi et al., 2023; A.b. et al., 2023; Zhou and Zafarani, 2020; Shu et al., 2017)), to the best of our knowledge, there is no survey on automatic credibility assessment and detection of credibility signals from the NLP perspective. In this paper, we therefore provide the first comprehensive and systematic literature overview of credibility signal detection with the focus on NLP and LLMs. The main contributions are as follows:
We address the fragmentation of existing works on automatic credibility assessment and credibility signals detection. We connect these research efforts under a common framework and identify shared trends, challenges, and open problems. In this way, we motivate future research in connecting and integrating individual outcomes (i.e., to move from currently isolated individual signals to credibility assessment effectively combining them).
@@
We provide a necessary background overview of definitions and dimensions of credibility assessment (including a unified categorization of credibility signals based on their various taxonomies and list used in the existing literature) to setup the common understanding that is currently missing in the existing works.
@@
We analyse and describe a total of 175 NLP works tackling the automatic credibility assessment and detection of credibility signals. We provide an in-depth description of 3 key categories of credibility signals, which have been selected due to a considerable NLP research interest and their recognition as important by end users, namely: (i) factuality, subjectivity and bias; (ii) persuasion techniques and logical fallacies; and (iii) check-worthy and fact-checked claims. This analysis is complemented with an overview of additional 6 categories of credibility signals. We also thoroughly position such NLP research in the context of other modalities as well as multidisciplinary works.
We identify and critically analyse key gaps in the existing literature, highlighting underexplored areas and methodological limitations in the current approaches to credibility assessment. Furthermore, we outline future research challenges and opportunities, with particular emphasis on the rapidly evolving landscape shaped by generative AI, especially Large Language Models (LLMs). We discuss how these advancements both pose new threats, such as the large-scale generation of highly credible yet false content, and offer new capabilities for enhancing the detection and aggregation of credibility signals. Our analysis aims to guide future research toward more comprehensive, robust and explainable credibility assessment approaches.
@@
Besides the review of published research works, this survey builds on extensive past research activities and acquired knowledge of its authors in the target area. A unique composition of authors consisting of experts not only on the primary NLP area, but also on other related research domains (such as computer vision, social network analysis) as well as media professionals, provides a novel holistic perspective on the addressed topic. Such diversity contributes, besides others, to identification of open problems and future challenges, in which an application of research in practice plays a prominent role.
@@
This survey is structured into 10 sections. Section 2 provides definitions of core concepts related to credibility, describes various dimensions of credibility assessment as well as defines the scope of this survey while situating it into the context of other existing surveys. In Section 3, the methodology employed in the survey process is thoroughly described. Section 4 focuses on research papers that address automatic credibility assessment (i.e., approaches addressing both steps — detection of credibility signals and their aggregation into a credibility label/score). Subsequently, Sections 5-7 address an automatic detection of 3 selected key categories of credibility signals. Section 8 supplements this analysis with a brief overview of additional 6 categories of credibility signals, including the complementary perspective of credibility signals addressed in the non-NLP research. In Section 9, we provide an orthogonal analysis of challenges and open problems that characterize the state of the art in this area. Finally, conclusions are drawn in Section 10.
@@ 2.1. Definitions
The construct of credibility has received more scholarly attention than most other communication variables, with foundational research dating back to 1951, when Hovland and Weiss (1951) demonstrated that the effectiveness of communication is heavily influenced by the perceived credibility of its source. A systematic meta-analysis of leading international communication journals spanning 1951 to 2011 (Hellmueller and Trilling, 2012) examined how source, message, and media credibility have been conceptualized and measured over time. The findings revealed significant inconsistencies across credibility scales, a lack of operational precision, and limited replication and validation of the credibility construct. These discrepancies largely stem from the fact that credibility is a complex, multidimensional concept that lacks a single, unified, and widely accepted definition.
Existing definitions vary especially across various disciplines and fields, which study credibility from different perspectives, such as information science, psychology, marketing or human-computer interaction (HCI) (Rieh and Danielson, 2007). It is commonly defined as a high-level construct with a help of related concepts, such as believability (which is considered to be roughly a synonym with credibility), trust/trusthworthiness, reliability or their various combinations (Qureshi et al., 2021; Alrubaian et al., 2019; Rieh and Danielson, 2007; Self and Roberts, 2019). Credibility is always tied to a target entity (an object of assessment), that can be either a piece of content or a source spreading such content (Kiousis, 2001). Due to ambiguity of terminology, in this survey we opt for community-contributed definitions (Hawke, [n. d.]) gathered by the Credible Web Community Group111https://credweb.org/. It defines credibility as a degree to which information is credible (believable) and to which information appears non-misleading and useful (for the given audience).
Credibility assessment is a process of ascertaining some level/degree of credibility to a target entity. Practically, credibility assessment can be formulated as (Qureshi et al., 2021): (i) an ordinal classification problem (i.e., assigning a label from pre-defined categories, such as a low/high level of credibility); or as (ii) a regression/scoring problem (i.e., assigning a numerical score). In both cases, there are no standard scales adopted yet. Credible Web Community Group distinguishes two kinds of credibility assessment acts (Hawke, [n. d.]):
Making credibility decisions is an act of adhoc deciding for oneself what to believe, which is often informal, immediate, and unconscious. Doing this incorrectly often leads to being misled, although it may not be practical to do it correctly at all times.
Credibility analysis is an act of systematic gathering, organizing, and analysing evidence to help people make credibility decisions about a particular information item. A credibility analysis process (performed manually by people or automatically by machines) might produce some kind of report which might itself be called a credibility assessment and might include a credibility score. In the following text, we will focus especially on this act.
@@
Credibility indicator (Hawke, [n. d.]) is commonly used interchangeably with credibility signal. In some communities, credibility signal is used for inputs to credibility assessment algorithms and credibility indicator is used for the display features added to the output, to communicate (explain) results to an end user. In this survey, we prefer to use credibility signal, while we would like to emphasize that in practice many signals can be used directly also as indicators communicated to end users.
Finally, credibility assessment tools (Hawke, [n. d.]) are software features or applications which perform credibility assessment or help people do so. Such tools can either derive credibility signals, or implement also a credibility assessment algorithm to aggregate such signals into a credibility label/score.
Credibility and credibility assessment is inherently very close to false information and the task of false information detection. Nevertheless, credibility (as also defined above) is rather a parallel/complementary concept. While they tend to correlate (low credibility may indicate false content), non-credible content does not necessarily need to be false only and vice versa.
We would like to emphasize that credibility signals have been already recognized in false information research as a possible solution towards more accurate and explainable methods. The work by Grieve and Woodfield (2023) delve into the complexities of identifying and analysing disinformation deceptive news. The authors study the way language adapts to different contexts and purposes, arguing that the linguistic choices in deceptive news differ systematically from those in genuine news. The difference in intent – to deceive versus to inform – leads to detectable variations in language use. The authors illustrate this with the famous case of fake news involving Jayson Blair of The New York Times, where false news were less informationally dense and less confident than real ones. The pressure to invent news and the lack of factual grounding influenced writing style, making it less precise and authoritative. The authors propose that this type of register variation could be a broader indicator of fake news. They call for continued research to understand the linguistic patterns of fake news and to develop strategies to combat its spread. Although the above framework is only based on the analysis of one reporter’s articles, this approach may be extended to corpora of non-credible/fake news and credible/real news on various topics to try to identify credibility signals/register variation between the two.
@@ 2.2. Dimensions of credibility assessment
In the past research works, a wide range of approaches studying the concept of credibility and credibility assessment was developed. In the following overview (see also Figure 1), we provide various dimensions adopted to study these concepts.
Approaches to credibility assessment. First, the credibility assessment methods can be broadly divided into three major groups (Alrubaian et al., 2019):
A large group of automation-based approaches focuses either on automatically detecting the credibility signals or aggregating them into the credibility indicators. The complexity of such automation can be very diverse, ranging from simple rule- and heuristic-based methods, through weighted and information retrieval (IR) approaches to supervised/unsupervised machine learning. This group also covers research works focusing on building datasets appropriate to create such methods/models; as well as systems (web applications, browser extensions, etc.) built on the top of such automated methods.
The second group corresponds to human-based approaches, in which various aspects of credibility are studied from the perspective of end-users. The human-based credibility assessment approaches can range from voting methods, cognitive and perception approaches to manual verification approaches.
Finally, hybrid approaches combine and utilize the advantages of both the automation- and human-based approaches.
@@
Content level. At the content level, the task is to analyse the content attributes typically of a social media post or a news article. It is the most fundamental and important type of assessment, which is commonly used to determine the credibility at other levels (e.g., if a post is assessed to be credible, also a corresponding user/source or an associated topic/event is considered to be credible).
@@
First, reflecting the source of information, many works broadly distinguish between content- and context-based credibility signals (Zhang et al., 2018; Kang et al., 2012; Bhuiyan et al., 2020; Wu et al., 2020; El Ballouli et al., 2017; Esteves et al., 2018; Qureshi and Malick, 2023; Piccolo et al., 2021; Jahanbakhsh and Karger, 2024; Chang et al., 2021; Ji et al., 2023b; Gao et al., 2015; Choi et al., 2023). Content-based signals are derived from the content itself, such as partisanship, emotional appeal, or persuasion techniques. On the other hand, context-based signals, also referred to as provenance- (Feng et al., 2023) or meta-information, take into account contextual (e.g., user, source) cues, such as user statistics, location, domain reputation, or the number of shares. In practice, when a content-level credibility (e.g., a credibility of a social media post) is about to be assessed, a combination of both – content-based credibility signals (e.g., a presence of persuasion techniques) as well as context-based credibility signals (e.g., the same claim was already fact-checked) – may be utilized.